# Feynman Formula for Discrete-time Quantum Walks

### v1

@@ 1. Introduction
There are two types of quantum walks, continuous-time and discrete-time quantum walks. In this paper, we only discuss discrete-time quantum walks which were introduced by Aharonov, Davidovich and Zagury [Aharonov et al., 1993] and began to draw attention to the quantum computing community since the paper [Ambainis et al., 2001]. We refer readers to the numerous surveys [Kempe, 2003], [Venegas-Andraca, 2012], [Konno, 2020], [Zhou, 2021], [Kadian et al., 2021], ordered from the earliest to the most recent, and the references therein for detailed development of the literature of quantum walks.
@@
Despite being defined to quantize classical random walks, quantum walks are deterministic dynamics and were also termed “quantum cellular automata” by [Meyer, 1996]; they are well-known to behave very differently from classical random walks. The characteristic example is its famous ballistic long-time behavior with a bimodal weak limit (see for example [Konno, 2005], [Grimmett et al., 2004]), which is in big contrast to the diffusive behavior driven by the classical central limit theorem. Thus, it is not surprising that techniques for studying quantum walks are rather non-probabilistic. They include, but are not limited to, path-counting and orthogonal polynomials [Konno, 2005], Fourier transform [Grimmett et al., 2004] [Maeda et al., 2018] [Strauch, 2006], the energy method [Maeda and Suzuki, 2020] and matrix analysis [Cantero et al., 2012] (see also the survey [Konno, 2020]).
@@
In addition, the ballistic behavior described in the previous paragraph is also reflected by the possibility of a hyperbolic space-time rescaling of the quantum walk to a system of transport PDEs with a phase interaction term. This was first stated and termed “quantum transport PDEs” in [Molfetta and Debbasch, 2012] (see also [Maeda et al., 2018] for discussions of this type of Dirac PDEs). We refer the reader to the recent book [Molfetta, 2024] and references therein for developments in Physics between quantum walks and transport PDEs.
@@
Our main contribution is to write a Feynman-type representation for quantum walks, bringing in probabilistic techniques for studying quantum walks. With our Feynman formula, we explicitly recognize the amplitude of a quantum walk to be a conditional expectation of a classical four-state Markov additive process. Using the formula, we are able to relate the spectral decomposition of the Markov additive process and the limiting density of quantum walks. In addition, we provide a probabilistic representation for the quantum transport PDE in terms of continuous-time scaling limits of the aforementioned four-state Markov additive process.
@@
Connections between classical processes and quantum walks have been widely studied in the literature. One direction is the problem of mimicking. Since quantum walks are unitary dynamics, taking modulus squares to the amplitudes of states in each step gives a flow of probability marginals if the walk is initialized with a normalized state. It has been studied in [Montero, 2017] and [Yamagami et al., 2022] how to match this flow of marginals with a time-inhomogeneous Markov chain. In addition, it is well-established in the continuous-time literature (see [Carlen, 1984]) that flow of marginals induced by a Schrödinger equation can be mimicked by a homogeneous continuous-time Markov process, called a conservative diffusion. In addition, to incorporate the effect of decoherence to quantum walks, [Attal et al., 2012] introduced open quantum walks; by repeatedly measuring and running an open quantum walk, one naturally obtains quantum trajectories, which are classical stochastic processes. See the survey article [Sinayskiy and Petruccione, 2019] and the introduction section in [Carbone et al., 2022] for details. Finally, by introducing a walk-type zeta function, [Konno and Sato, 2012], [Konno et al., 2020], [Komatsu et al., 2022] recognized universal principles underlying both quantum walks and classical walks with Markov increments so that the walks can be handled in a unifying scheme. We remark that the 4-state Markov additive process induced from our Feynman formulae does not come from any measurement procedure and we intend to use it as a technique to uncover properties of quantum walks instead of comparing with quantum walks under a unifying scheme. Therefore, we believe that the relation between classical processes and quantum walks that we propose is novel and useful.
@@
The paper is structured as follows. In Section 2, we recall the definition of a one-dimensional quantum walk, derive our main Feynman formula for the rotation coin $e^{i\theta\sigma_{1}}$, and comment on potential Monte Carlo simulations. Next, in Section 3, we derive a relation between the spectral decomposition of the Markov additive process and the limiting density of the homogeneous quantum walk , and we demonstrate the derivation of the weak limit for the long-term behavior of quantum walks. In Section 4, we consider a space/time rescaling of quantum walks, which leads to a system of transport PDEs in continuous time and space, with a phase interaction term. Monte Carlo simulations are very efficient to approximate the solution of such a system. Finally, in Section 5, we propose extensions to our Feynman formula to more general quantum walks, which include more general homogeneous coins, time-dependent coins (see for example [Bañuls et al., 2006]) and site-dependent coins (see for example [Wójcik et al., 2004], [Linden and Sharam, 2009], [Maeda et al., 2018] and [Maeda and Suzuki, 2020]).
@@ 2.1. Notations
For common notations in the quantum literature, see [Nielsen and Chuang, 2010] and [Hall, 2013] for a more quantum computational and quantum mechanical account respectively. Here we mainly follow notations in [Watrous, 2018] with slight modifications.
@@ 2.2. One-dimensional quantum walks
Quantum walks on $\mathbb{Z}$ are discrete-times dynamics of the form
(2.3) $\psi_{n+1}=W\psi_{n},\quad\psi_{0}\in L(\ell_{2}(\mathbb{Z})\otimes\ell_{2}(\{\pm 1\}))$
@@
where $C_{x}\in U(2)$ is the coin at position $x\in\mathbb{Z}$. Denote the action of the coin by $C_{x}\begin{pmatrix}\alpha_{x}\\ \beta_{x}\end{pmatrix}:=\begin{pmatrix}\tilde{\alpha}_{x}\\ \tilde{\beta}_{x}\end{pmatrix}$.
Then, the shift step gives:
@@
With the homogeneous balanced coin $C_{x}\equiv e^{i(\pi/4)\sigma_{1}}=\frac{1}{\sqrt{2}}\begin{bmatrix}1&i\\ i&1\end{bmatrix}$ for all $x\in\mathbb{Z}$ and the normalized initial condition $\psi_{0}(0,-1)=\psi_{0}(0,1)=1/\sqrt{2}$, $\psi_{0}(x,\pm 1)\equiv 0$ for all $x\neq 0$, the first two steps of this quantum walk is represented in Figure 1.
@@
Figure 1. The mechanism of the one-dimensional quantum walk $(\psi_{n}=W^{n}\psi_{0},n=0,1,2)$ with the balanced coin $e^{i(\pi/4)\sigma_{1}}$ and the initial condition $\psi_{0}(0,-1)=\psi_{0}(0,1)=1/\sqrt{2}$ in Example 2.2. Two coin states z=1z=-1 and $z=+1$ are assigned for each position $x\in\mathbb{Z}$. For each step, we apply the coin operator $C$ first and then the shift operator $S$. At $n=2$, the state $C\psi_{1}$ has 4 non-zero components because the action of $C$ splits each component of $\psi_{1}$ into two.
@@ 2.3. Deriving the Feynman formula
Introducing a sequence of i.i.d. Poisson random variables $(N_{1},N_{2},\cdots)$ with parameter $\theta\in(0,2\pi)$, we define the Markov additive model [^1] A Markov additive model is a triple that consists of a Markov chain, a real-valued function of it and a sum of that real-valued function of the Markov chain (see for instance [Ney and Nummelin, 1987a] and [Ney and Nummelin, 1987b]). $(S_{n},Z_{n}:=f(S_{n}),X_{n})\in\{0,1,2,3\}\times\{\pm 1\}\times\mathbb{Z}$ by
@@ 2.4. Monte Carlo simulations
where $(S^{(k)}_{n},Z^{(k)}_{n},X^{(k)}_{n})$ are computed from $(N_{1}^{(k)},\cdots,N_{n}^{(k)})$ according to the formulas in Definition 2.6. Well, we tried and that doesn’t work well unless using a huge unrealistic number of realizations even for a relatively small time $n$. The detailed study of the sample size to compute it with stochastic simulation methods that include importance sampling, e.g., [Chatterjee and Diaconis, 2018] is an ongoing project. In fact, we are in a comparable situation with the stochastic model for the telegrapher’s equation described by the great Mark Kac in [Kac, 1974]. Here is what he wrote about implementing a Monte Carlo computation for a discrete time transport system:
@@
However, in this same article, Mark Kac suggests a rescaling of the system to pass to the continuous-time limit. He obtained a probabilistic representation of the solution of a transport system of Boltzmann type and observed that in this limiting situation, Monte Carlo technique makes perfect sense and is amazingly efficient. This has been the basis and widely used technique for solving transport equations with Boltzmann interactions (see [Cessenat et al., 1989]).
@@ 3.1. Motivation - ballistic weak limit of quantum walk
It first appeared in Konno’s work [Konno, 2005] that if we start with the state
@@
which could be found in, for instance, [Konno, 2005, Sunada and Tate, 2012] with slightly different notations. Indeed one can check that the function $f_{K}\cdot\omega_{\psi_{0}}$ is non-negative and integrates to one over $\mathbb{R}$. In order to derive this convergence, we characterize the distribution of ${\Xi_{n}}/{n}$ by
@@
Rescaling the space in (3.5) by setting $x=an$ for $a\in\mathbb{R}$ such that $an\in\mathbb{Z}$, by the theory of large deviations [Dembo and Zeitouni, 2010], we see that each probability term $\mathbb{P}(T_{n}/n=az,S_{n}=\cdot)$ in (3.5) is exponentially small for large $n$ and $a\neq 0$. Equation (3.5) tells us that the quantum walk amplitude $\psi_{n}^{\xi}(an,z)$ is a re-scaled (by a factor $e^{n\theta}$) difference of small probabilities led by large deviation.
@@ 3.3.1. Quantum walk amplitudes
where $\operatorname{diag}(e^{ik},e^{-ik})e^{i\theta\sigma_{1}}$ is the matrix representation of the quantum walk operator $W:=SC$ over the Fourier domain in space, which is first documented in [Grimmett et al., 2004]. One can easily show that (3.16) in Theorem 3.3 still holds using this matrix representation, which coincides with the integral studied in Equation (1.20) in [Sunada and Tate, 2012]. In the following, we continue with our initial natural choice of basis in Proposition 3.2.
@@ 3.3.2. Deriving quantum walks asymptotics
This is a consequence of the stationary phase asymptotics uniformly in $a\in[y_{1},y_{2}]$. Note that for fixed $\epsilon>0$, we can choose $n$ large enough such that the inequality $\cos\theta+\epsilon/2\leq\frac{1}{n}\left\lfloor ny_{1}\right\rfloor\leq\frac{1}{n}\left\lfloor ny_{2}\right\rfloor\leq\cos\theta-\epsilon/2$ holds. See Appendix A for the details. ∎
@@
4. Another scaling limit and relation with the telegraph process
@@
It is now very natural to write a system of quantum transport PDEs solved by $\psi$, for which we have the probabilistic representation (4.5). In fact, this type of PDEs, known as Dirac PDEs, has been studied in the Physics literature and we refer to the papers [Maeda et al., 2018] and [Maeda and Suzuki, 2020] where nonlinear PDEs are also discussed.
@@
Let $h>0$ and denoting $\mathbb{E}_{0,y,z}$ the expectation over the initial condition $(N_{0},Y_{0},Z_{0})=(0,y,z)$, We have
(4.10) $\qquad\psi(t+h,y,z)=e^{(t+h)\lambda}\mathbb{E}_{0,y,z}\Big[i^{N_{t+h}}f\Big(y-z\int_{0}^{t+h}(-1)^{N_{s}}\mathrm{d}s,z(-1)^{N_{t+h}}\Big)\Big]$
@@
$=\mathbb{E}\left[i^{M_{t}}f\Big(Y_{h}-Z_{h}\int_{h}^{t+h}(-1)^{M_{s-h}}\mathrm{d}s,Z_{h}(-1)^{M_{s}}\Big)\right]$
$=e^{-t\lambda}\psi(t,Y_{h},Z_{h}).$
@@
with $u(0,y,z)=f(y,z)$ and where $u_{\pm}(t,y):=u(t,y,\pm 1)$. This system of equations is similar to a linear Boltzmann transport system (see [Kac, 1974]), except that the interaction term is now a phase interaction term; it is termed a system of ”quantum transport equations” in [Molfetta and Debbasch, 2012]. Our novelty here is to provide a probabilistaic formulation of the solution to the PDE.

### v2

@@ 1. Introduction
There are two types of quantum walks, continuous-time and discrete-time quantum walks. In this paper, we only discuss discrete-time quantum walks which were introduced by Aharonov et al., [1993] and began to draw attention to the quantum computing community since the paper Ambainis et al., [2001]. We refer readers to the numerous surveys Kempe, [2003], Venegas-Andraca, [2012], Konno, [2020], Zhou, [2021], Kadian et al., [2021], ordered from the earliest to the most recent, and the references therein for detailed development of the literature of quantum walks.
@@
Despite being defined to quantize classical random walks, quantum walks are deterministic dynamics and were also termed “quantum cellular automata” by Meyer, [1996]; they are well-known to behave very differently from classical random walks. The characteristic example is its famous ballistic long-time behavior with a bimodal weak limit (see for example Konno, [2005], Grimmett et al., [2004]), which is in big contrast to the diffusive behavior driven by the classical central limit theorem. Thus, it is not surprising that techniques for studying quantum walks are rather non-probabilistic. They include, but are not limited to, path-counting and orthogonal polynomials Konno, [2005], Fourier transform Grimmett et al., [2004] Maeda et al., [2018] Strauch, [2006], the energy method Maeda and Suzuki, [2020] and matrix analysis Cantero et al., [2012] (see also the survey Konno, [2020]).
@@
In addition, the ballistic behavior described in the previous paragraph is also reflected by the possibility of a hyperbolic space-time rescaling of the quantum walk to a system of transport PDEs with a phase interaction term. This was first stated and termed “quantum transport PDEs” in Molfetta and Debbasch, [2012] (see also Maeda et al., [2018] for discussions of this type of Dirac PDEs). We refer the reader to the recent book Molfetta, [2024] and references therein for developments in Physics between quantum walks and transport PDEs.
@@
Our main contribution is to write a Feynman-type representation for quantum walks, bringing in probabilistic techniques for studying them. With our Feynman formula, we explicitly recognize the amplitude of a quantum walk to be a conditional expectation of a classical four-state Markov additive process. Using the formula, we are able to relate the spectral decomposition of the Markov additive process and the limiting density of quantum walks. In addition, we provide a probabilistic representation for the quantum transport PDE in terms of continuous-time scaling limits of the aforementioned four-state Markov additive process. This offers an efficient Monte Carlo method for such PDE.
@@
Connections between classical processes and quantum walks have been widely studied in the literature. One direction is the problem of mimicking. Since quantum walks are unitary dynamics, taking modulus squares to the amplitudes of states in each step gives a flow of probability marginals if the walk is initialized with a normalized state. It has been studied in Montero, [2017] and Yamagami et al., [2022] how to match this flow of marginals with a time-inhomogeneous Markov chain. In addition, it is well-established in the continuous-time literature (see Carlen, [1984]) that flow of marginals induced by a Schrödinger equation can be mimicked by a homogeneous continuous-time Markov process, called a conservative diffusion. In addition, to incorporate the effect of decoherence to quantum walks, Attal et al., [2012] introduced open quantum walks; by repeatedly measuring and running an open quantum walk, one naturally obtains quantum trajectories, which are classical stochastic processes. See the survey article Sinayskiy and Petruccione, [2019] and the introduction section in Carbone et al., [2022] for details. Finally, by introducing a walk-type zeta function, Konno and Sato, [2012], Konno et al., [2023], Komatsu et al., [2022] recognized universal principles underlying both quantum walks and classical walks with Markov increments so that the walks can be handled in a unifying scheme. We remark that the 4-state Markov additive process induced from our Feynman formulae does not come from any measurement procedure and we intend to use it as a technique to uncover properties of quantum walks instead of comparing with quantum walks under a unifying scheme. Therefore, we believe that the relation between classical processes and quantum walks that we propose is novel and useful.
@@
The paper is structured as follows. In Section 2, we recall the definition of a one-dimensional quantum walk, derive our main Feynman formula for the rotation coin $e^{i\theta\sigma_{1}}$, and comment on potential Monte Carlo simulations. Next, in Section 3, we derive a relation between the spectral decomposition of the Markov additive process and the limiting density of the homogeneous quantum walk , and we demonstrate the derivation of the weak limit for the long-term behavior of quantum walks. In Section 4, we consider a space-time rescaling of quantum walks, which leads to a system of transport PDEs in continuous time and space, with a phase interaction term. Monte Carlo simulations are very efficient to approximate the solution of such a system. Finally, in Section 5, we propose extensions to our Feynman formula to more general quantum walks, which include more general homogeneous coins, time-dependent coins (see for example Bañuls et al., [2006]) and site-dependent coins (see for example Wójcik et al., [2004], Linden and Sharam, [2009], Maeda et al., [2018] and Maeda and Suzuki, [2020]).
@@ 2.1. Notations
For common notations in the quantum literature, see Nielsen and Chuang, [2010] and Hall, [2013] for a more quantum computational and quantum mechanical account respectively. Here we mainly follow notations in Watrous, [2018] with slight modifications.
@@ 2.2. One-dimensional quantum walks
Quantum walks on $\mathbb{Z}$ are discrete-time dynamics of the form
(2.3) $\psi_{n}=W\psi_{n-1}\quad\mbox{for}\quad n\geq 1;\quad\psi_{0}\in L(\ell_{2}(\mathbb{Z})\otimes\ell_{2}(\{\pm 1\})),$
@@
where $C_{x}\in U(2)$ is the coin at position $x\in\mathbb{Z}$.
If we denote the action of the coin by $C_{x}\begin{pmatrix}\alpha_{x}\\ \beta_{x}\end{pmatrix}:=\begin{pmatrix}\tilde{\alpha}_{x}\\ \tilde{\beta}_{x}\end{pmatrix}$, then, the shift step gives:
@@
With the homogeneous rotation coin $C_{x}\equiv e^{i(\pi/4)\sigma_{1}}=\frac{1}{\sqrt{2}}\begin{bmatrix}1&i\\ i&1\end{bmatrix}$ for all $x\in\mathbb{Z}$ and the normalized initial condition $\psi_{0}(0,-1)=\psi_{0}(0,1)=1/\sqrt{2}$, $\psi_{0}(x,\pm 1)\equiv 0$ for all $x\neq 0$, the first two steps of this quantum walk is represented in Figure 1.
@@
Figure 1. The mechanism of the one-dimensional quantum walk $(\psi_{n}=W^{n}\psi_{0},n=0,1,2)$ with the rotation coin $e^{i(\pi/4)\sigma_{1}}$ and the initial condition $\psi_{0}(0,-1)=\psi_{0}(0,1)=1/\sqrt{2}$ in Example 2.2. Two coin states z=1z=-1 and $z=+1$ are assigned for each position $x\in\mathbb{Z}$. For each step, we apply the coin operator $C$ first and then the shift operator $S$. At $n=2$, the state $C\psi_{1}$ has 4 non-zero components because the action of $C$ splits each component of $\psi_{1}$ into two.
@@ 2.3. Deriving the Feynman formula
Introducing a sequence of i.i.d. Poisson random variables $(N_{1},N_{2},\cdots)$ with parameter $\theta\in(0,2\pi)$, we define the Markov additive model [^1] A Markov additive model is a triple that consists of a Markov chain, a real-valued function of it and a sum of that real-valued function of the Markov chain (see for instance Ney and Nummelin, 1987a and Ney and Nummelin, 1987b ). $(S_{n},Z_{n}:=f(S_{n}),X_{n})\in\{0,1,2,3\}\times\{\pm 1\}\times\mathbb{Z}$ by
@@ 2.4. Monte Carlo simulations
where $(S^{(k)}_{n},Z^{(k)}_{n},X^{(k)}_{n})$ are computed from $(N_{1}^{(k)},\cdots,N_{n}^{(k)})$ according to the formulas in Definition 2.6. Well, we tried and that doesn’t work well unless using a huge unrealistic number of realizations even for a relatively small time $n$. The detailed study of the sample size to compute it with stochastic simulation methods that include importance sampling, e.g., Chatterjee and Diaconis, [2018] is an ongoing project. In fact, we are in a comparable situation with the stochastic model for the telegrapher’s equation described by the great Mark Kac in Kac, [1974]. Here is what he wrote about implementing a Monte Carlo computation for a discrete time transport system:
@@
However, in this same article, Mark Kac suggests a rescaling of the system to pass to the continuous-time limit. He obtained a probabilistic representation of the solution of a transport system of Boltzmann type and observed that in this limiting situation, Monte Carlo technique makes perfect sense and is amazingly efficient. This has been the basis and widely used technique for solving transport equations with Boltzmann interactions (see Cessenat et al., [1989]).
@@ 3.1. Motivation - ballistic weak limit of quantum walk
It first appeared in Konno’s work Konno, [2005] that if we start with the state
@@
which could be found in, for instance, Konno, [2005]; Sunada and Tate, [2012] with slightly different notations. Indeed one can check that the function $f_{K}\cdot\omega_{\psi_{0}}$ is non-negative and integrates to one over $\mathbb{R}$. In order to derive this convergence, we characterize the distribution of ${\Xi_{n}}/{n}$ by
@@
Rescaling the space in (3.5) by setting $x=an$ for $a\in\mathbb{R}$ such that $an\in\mathbb{Z}$, by the theory of large deviations Dembo and Zeitouni, [2010], we see that each probability term $\mathbb{P}(T_{n}/n=az,S_{n}=\cdot)$ in (3.5) is exponentially small for large $n$ and $a\neq 0$. Equation (3.5) tells us that the quantum walk amplitude $\psi_{n}^{\xi}(an,z)$ is a re-scaled (by a factor $e^{n\theta}$) difference of small probabilities led by large deviation.
@@ 3.3.1. Quantum walk amplitudes
where $\operatorname{diag}(e^{ik},e^{-ik})e^{i\theta\sigma_{1}}$ is the matrix representation of the quantum walk operator $W:=SC$ over the Fourier domain in space, which is first documented in Grimmett et al., [2004]. One can easily show that (3.16) in Theorem 3.3 still holds using this matrix representation, which coincides with the integral studied in Equation (1.20) in Sunada and Tate, [2012]. In the following, we continue with our initial natural choice of basis in Proposition 3.2.
@@ 3.3.2. Deriving quantum walks asymptotics
This is a consequence of the stationary phase asymptotics uniformly in $a\in[y_{1},y_{2}]$. Note that for fixed $\epsilon>0$, we can choose $n$ large enough such that the inequality $\cos\theta+\epsilon/2\leq\frac{1}{n}\left\lfloor ny_{1}\right\rfloor\leq\frac{1}{n}\left\lfloor ny_{2}\right\rfloor\leq\cos\theta-\epsilon/2$ holds. See Appendix for the details. ∎
@@
4. Hyperbolic scaling limit and relation with the telegraph process
@@
It is now very natural to write a system of quantum transport PDEs solved by $\psi$, for which we have the probabilistic representation (4.5). In fact, this type of PDEs, known as Dirac PDEs, has been studied in the Physics literature and we refer to the papers Maeda et al., [2018] and Maeda and Suzuki, [2020] where nonlinear PDEs are also discussed.
@@
Let $h>0$ and denoting by $\mathbb{E}_{0,y,z}$ the expectation over the initial condition $(N_{0},Y_{0},Z_{0})=(0,y,z)$, we have
(4.10) $\qquad\qquad\psi(t+h,y,z)=e^{(t+h)\lambda}\mathbb{E}_{0,y,z}\Big[i^{N_{t+h}}f\Big(y-z\int_{0}^{t+h}(-1)^{N_{s}}\mathrm{d}s,z(-1)^{N_{t+h}}\Big)\Big]$
@@
$=\mathbb{E}\left[i^{M_{t}}f\Big(Y_{h}-Z_{h}\int_{h}^{t+h}(-1)^{M_{s-h}}\mathrm{d}s,Z_{h}(-1)^{M_{s}}\Big)\right]=e^{-t\lambda}\psi(t,Y_{h},Z_{h}).$
@@
with $u(0,y,z)=f(y,z)$ and where $u_{\pm}(t,y):=u(t,y,\pm 1)$. This system of equations is similar to a linear Boltzmann transport system (see Kac, [1974]), except that the interaction term is now a phase interaction term; it is termed a system of ”quantum transport equations” in Molfetta and Debbasch, [2012]. Our novelty here is to provide a probabilistaic formulation of the solution to the PDE.